The length of a rectangle is 5 m more than its width. What is the width of the rectangle if it's area is 84 m^2?
Answer by jorel555(1290)

l=w+5
w(w+5)=84
w^2+5w-84=0
(w+12)(w-7)=84
w=-12,7
The width is 7m, the length is 12m!!!!!!!!!!!!!!